# Quiver moduli and quantum loop algebras

### v1

@@ Abstract.
We explore the relationship between representations of quantum loop algebras, shuffle algebras, and moduli of quiver representations, with the goal of proving an equality of numbers arising from these three points of view.
@@ 1.1. The setting
Let ${\mathbb{K}}={\mathbb{Q}}(q)$ and fix a Kac-Moody Lie algebra $\mathfrak{g}$}, with a set $I$ of simple roots. We write $(d_{ij})_{i,j\in I}$ for the symmetrized generalized Cartan matrix associated to $\mathfrak{g}$}, see (60). The representation theory of the quantum loop algebra
@@
has long been studied, see for instance [6, 24] for $\mathfrak{g}$} of finite type and [37] for $\mathfrak{g}$} of general type. We will consider the simple modules $L({\boldsymbol{\psi}})$ of the Borel subalgebra of $U_{q}(L{\mathfrak{g}})$ which are indexed by so-called highest $\ell$-weights
@@ 1.2. Shuffle algebras
${\mathcal{S}}=\bigoplus_{{\boldsymbol{n}}\in{\mathbb{N}^{I}}}{\mathcal{S}}_{{\boldsymbol{n}}}\subseteq\bigoplus_{{\boldsymbol{n}}\in{\mathbb{N}^{I}}}{\mathcal{V}}_{{\boldsymbol{n}}}={\mathcal{V}},\text{ where }{\mathcal{V}}_{{\boldsymbol{n}}}={\mathbb{K}}[z_{i1}^{\pm 1},\dots,z_{in_{i}}^{\pm 1}]^{\textrm{sym}}_{i\in I}$
@@
(3) $L({\boldsymbol{\psi}})=L^{\neq 0}({\boldsymbol{\psi}})\otimes L^{\text{ord }{\boldsymbol{\psi}}}$
where the graded vector space $L^{\text{ord }{\boldsymbol{\psi}}}$ only depends on the orders of the rational functions $\psi_{i}(z)$ at $z=0$, and was described in [38]. Meanwhile, the vector space $L^{\neq 0}({\boldsymbol{\psi}})$ can be shown in [26] to match the underlying graded vector space of the same-named simple module of an appropriately shifted quantum loop algebra. We proved in [37] that this vector space enjoys the following shuffle algebra description
@@
where $\Theta({\boldsymbol{\psi}})\subseteq{\mathcal{V}}$ is an explicit subset that we recall in Subsection 3.6. Moreover, the objects above admit gradings that will be denoted by
@@
see [38, Subsection 4.8]. Formula (3) implies the equality
$\chi_{q}(L({\boldsymbol{\psi}}))=\chi_{q}(L^{\neq 0}({\boldsymbol{\psi}}))\cdot\chi^{\text{ord }{\boldsymbol{\psi}}}$
@@ 1.3. Euler characteristics
The main purpose of the present paper is to propose a geometric realization of ${\mathcal{S}}_{{\boldsymbol{n}}}/\left({\mathcal{S}}_{{\boldsymbol{n}}}\cap\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}\right)$, inspired by the following seminal conjecture of Hernandez-Leclerc. We assume that  [^1] As has long been known to experts, working with $\boldsymbol{\psi}$} of the form (7) does not represent any actual restriction, since the $q$-character of an arbitrary $L({\boldsymbol{\psi}})$ naturally splits as a product of $q$-characters of $L({\boldsymbol{\psi}})$’s of the form (7): in other words, poles and zeroes of $\boldsymbol{\psi}$} that are in different geometric progressions with ratio $q$ do not interact with each other.
@@
Conjecture 1.4.
([25]) For $\mathfrak{g}$} of finite type and “reachable” highest $\ell$-weights $\boldsymbol{\psi}$},
@@
Conjecture 1.4 was proved in [29] by relating the $q$-characters of $L({\boldsymbol{\psi}})$ with a certain cluster algebra associated to the quiver $Q^{\text{gr}}$ (for which the relation to Euler characteristics of quiver moduli had been established on quite general grounds in [9]). This explains the adjective “reachable” in the statement of Conjecture 1.4: the proof of formula (8) only applies to those highest $\ell$-weights which correspond to cluster monomials under the correspondence developed in [29].
1.5. Toward general highest $\ell$-weights $\boldsymbol{\psi}$}
Although it might be excessive to hope that (8) holds as stated for all highest $\ell$-weights $\boldsymbol{\psi}$} of the form (7), it is quite reasonable to generalize it as follows: find a geometric formula for the coefficients of $\chi_{q}(L^{\neq 0}({\boldsymbol{\psi}}))$, in terms of the algebraic variety $N^{\text{stab}}_{\boldsymbol{x},{\boldsymbol{m}},{\mathbf{l}}}$. Since we already know from (6) that the aforementioned coefficients are given by the dimensions of the vector spaces ${\mathcal{S}}_{{\boldsymbol{n}}}/({\mathcal{S}}_{{\boldsymbol{n}}}\cap\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}})$, then the problem becomes equivalent to the following:
@@ 1.5. Toward general highest $\ell$-weights $\boldsymbol{\psi}$}
To set up our approach to the problem above, in what follows we fix arbitrary
@@
For any Kac-Moody Lie algebra $\mathfrak{g}$}, we have a ring isomorphism
@@
for any ${\boldsymbol{n}},\boldsymbol{x},{\boldsymbol{m}}$ as above, where ${\boldsymbol{\psi}}^{\emph{den}}$ denotes the denominator of (7). We write $K(X)$ for the 0-th topological $K$-theory of $X$. The space $M_{\boldsymbol{x},{\boldsymbol{m}}}^{\emph{stab}}$ has no odd $K$-theory.
@@
(15) ${\mathcal{V}}_{{\boldsymbol{n}}}\Big/\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}\xrightarrow{\sim}(\Delta_{{\mathbf{l}}}\cdot{\mathcal{V}}_{{\boldsymbol{n}}})\Big/\left((\Delta_{{\mathbf{l}}}\cdot{\mathcal{V}}_{{\boldsymbol{n}}})\cap\Theta\left(\frac{1}{{\boldsymbol{\psi}}^{\text{den}}}\right)_{\boldsymbol{x}}\right)\subset{\mathcal{V}}_{{\boldsymbol{n}}}\Big/\Theta\left(\frac{1}{{\boldsymbol{\psi}}^{\text{den}}}\right)_{\boldsymbol{x}}$
@@
For example, it is quite conceivable that for the reachable highest $\ell$-weights $\boldsymbol{\psi}$} to which Conjecture 1.4 applies, the direct image map
@@
is injective and its image coincides with the vector space that appears on the first line of (17). If $N_{\boldsymbol{x},{\boldsymbol{m}},{\mathbf{l}}}^{\text{stab}}$ could also be shown to not have any odd $K$-theory, then the previous sentence would provide an alternate proof of Conjecture 1.4.
@@ 1.7. Critical $K$-theory
Let us consider critical $K$-theory, which was developed in [1, 11, 27] (and more recently [4]) and applied to the understanding of simple modules $L({\boldsymbol{\psi}})$ in [45, 46]; see also [41] for the original definition of $K$-theoretic Hall algebras. We will start from a certain quiver with potential which has been studied in [21]: let $Q$ have vertex set $I$ and a single arrow $_{j}\square_{i}:i\rightarrow j$ for all $i,j\in I$ (this includes a single loop at every vertex $i$). Then we consider the potential
@@
where the spaces in the bottom row are analogues of those in the top row for the quiver of Conjecture 1.4, intersected with the stable open locus (see Subsection 4.6 for more details). In the bottom row, $\Delta_{{\mathbf{l}}}^{\text{gr}}$ is (a slightly renormalized version of) the product of those linear factors of (14) which vanish at $z_{ia}=x_{ia}$, specifically
@@
For $\mathfrak{g}$} of finite type and any highest $\ell$-weight $\boldsymbol{\psi}$}, the image of the composition $\rightarrow\rightarrow\downarrow$ in diagram (20) is naturally isomorphic to the vector space
@@
Conjecture 1.9.
The maps $\pi$ and $\iota^{\emph{gr},\emph{stab}}_{\boldsymbol{x},{\boldsymbol{m}}}$ in (20) are surjective and injective, respectively.
Combining Theorem 1.8 with Conjecture 1.9 and statement (22) implies that for $\mathfrak{g}$} of finite type, we have an isomorphism
We interpret (24) as the answer to problem (17), for $\mathfrak{g}$} of finite type. Indeed, the critical $K$-theory in the RHS of (24) is a replacement of the $K$-theory of the critical locus of $W^{\text{gr}}$, while multiplication by $\Delta^{\text{gr}}_{{\mathbf{l}}}$ is a derived version of imposing the vanishing of the $l_{i,s}$ linear combinations of framed paths that cut out the closed subvariety (9). Thus, the right-hand side of (24) is a replacement of $K(N_{\boldsymbol{x},{\boldsymbol{m}},{\mathbf{l}}}^{\text{stab}})\otimes_{{\mathbb{Q}}}{\mathbb{K}}$.
To conclude the introduction, let us compare the approach above to the earlier works [45, 46]: in loc. cit., the authors defined a convolution product on
@@
which they related to $U_{q}(L{\mathfrak{g}})$. Among other things, they defined prefundamental modules of $U_{q}(L{\mathfrak{g}})$ (i.e. the case ${\mathbf{l}}={\boldsymbol{0}}$ of (7)) using similar spaces as us. Our point of view is complementary to that of loc. cit.: we start from the vector space
@@
which we already know yields the simple module $L({\boldsymbol{\psi}})$, and then seek geometric realizations for the direct summands in (26). Once Conjecture 1.9 is proved, our point of view would provide not only a generalization of the Hernandez-Leclerc conjecture 1.4, but also a computation of critical convolution algebra modules (in the finite type case) in terms of shuffle algebras.
@@ 1.10. Acknowledgements
I would like to thank Ben Davison, David Hernandez, Shivang Jindal, Hiraku Nakajima, Tudor Pădurariu, Markus Reineke and Yukinobu Toda for numerous important conversations on the topics of the present paper.
@@ 2. Moduli of quiver representations
We recall the cell decomposition of moduli spaces of stable framed quiver representations from [12], and use it to describe the ring structure on the $K$-theory of the aforementioned moduli spaces (in a different way from the earlier work [17]). Hereafter, we write $K(X)$ for the 0-th topological $K$-theory ring of a smooth complex algebraic variety $X$, with rational coefficients. We thank Markus Reineke for sharing with us much advice pertaining to the contents of the present Section.

### v2

@@ Abstract.
A classic result of Hernandez-Leclerc and Kashiwara-Kim-Oh-Park relates the $q$-characters of so-called reachable simple modules of quantum affine algebras to the Euler characteristics of certain quiver moduli spaces. We categorify and generalize this relation to all simple modules in the Hernandez-Jimbo category $\mathcal{O}$} using critical $K$-theory (our results hold for quantum toroidal as well as quantum affine algebras).
@@ 1.1. The setting
Let ${\mathbb{K}}={\mathbb{Q}}(q)$ and fix a symmetrizable Kac-Moody Lie algebra $\mathfrak{g}$}, with a set $I$ of simple roots. We write $(d_{ij})_{i,j\in I}$ for the symmetrized generalized Cartan matrix of $\mathfrak{g}$}, see (63). The representation theory of the quantum loop algebra
@@
has long been studied, see for instance [8, 25] for $\mathfrak{g}$} of finite type and [38] for $\mathfrak{g}$} of general type. We will consider the simple modules $L({\boldsymbol{\psi}})$ of the Borel subalgebra of $U_{q}(L{\mathfrak{g}})$ (defined in loc. cit.) which are indexed by so-called highest $\ell$-weights
@@ 1.2. Shuffle algebras
${\mathcal{S}}=\bigoplus_{{\boldsymbol{n}}\in{\mathbb{N}^{I}}}{\mathcal{S}}_{{\boldsymbol{n}}}\subseteq\bigoplus_{{\boldsymbol{n}}\in{\mathbb{N}^{I}}}{\mathcal{V}}_{{\boldsymbol{n}}}={\mathcal{V}},\quad\text{where }{\mathcal{V}}_{{\boldsymbol{n}}}={\mathbb{K}}[z_{i1}^{\pm 1},\dots,z_{in_{i}}^{\pm 1}]^{\textrm{sym}}_{i\in I}$
@@
(3) $L({\boldsymbol{\psi}})\cong L^{\text{ord }{\boldsymbol{\psi}}}\otimes L^{\neq 0}({\boldsymbol{\psi}})$
where the graded vector space $L^{\text{ord }{\boldsymbol{\psi}}}$ only depends on the orders of the rational functions $\psi_{i}(z)$ at $z=0$, and was described in [39]. Meanwhile, the vector space $L^{\neq 0}({\boldsymbol{\psi}})$ will be shown in [27] to match the underlying graded vector space of the same-named simple module ([24]) of an appropriately shifted quantum loop algebra. We proved in [38] that this vector space has the following shuffle algebra description
@@
where $\Theta({\boldsymbol{\psi}})\subseteq{\mathcal{V}}$ is an explicit set of Laurent polynomials that we recall in Subsection 3.6. Moreover, the objects above admit gradings that will be denoted by
@@
see [39, Subsection 4.8]. Formula (3) implies the equality
$\chi_{q}(L({\boldsymbol{\psi}}))=\chi^{\text{ord }{\boldsymbol{\psi}}}\cdot\chi_{q}(L^{\neq 0}({\boldsymbol{\psi}}))$
@@ 1.3. Euler characteristics
The main purpose of the present paper is to produce a geometric realization of ${\mathcal{S}}_{{\boldsymbol{n}}}/\left({\mathcal{S}}_{{\boldsymbol{n}}}\cap\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}\right)$, inspired by the following seminal construction of Hernandez-Leclerc. We assume that  [^1] As has long been known to experts, working with $\boldsymbol{\psi}$} of the form (7) does not represent any actual restriction, since the $q$-character of an arbitrary $L({\boldsymbol{\psi}})$ naturally splits as a product of $q$-characters of $L({\boldsymbol{\psi}})$’s of the form (7): in other words, poles and zeroes of $\boldsymbol{\psi}$} that are in different geometric progressions with ratio $q$ do not interact with each other.
@@
Theorem 1.4.
(general conjecture in [26], proof for Kirillov-Reshetikhin $\boldsymbol{\psi}$} in loc. cit., general proof in [30]) For $\mathfrak{g}$} of finite type and “reachable” highest $\ell$-weights $\boldsymbol{\psi}$},
@@
Kashiwara-Kim-Oh-Park proved the result above by establishing a monoidal categorification of the cluster algebra associated to the quiver $Q^{\text{gr}}$ (for which the relation to Euler characteristics of quiver moduli had been established on quite general grounds in [10]), following the expectation of Hernandez-Leclerc. This explains the need for the adjective “reachable” in the statement of Theprem 1.4, since these are precisely the modules whose $q$-characters correspond to cluster monomials.
1.5. General highest $\ell$-weights $\boldsymbol{\psi}$}
We wish to generalize (8) to all highest $\ell$-weights $\boldsymbol{\psi}$}, even those which produce simple modules only of the Borel subalgebra of $U_{q}(L{\mathfrak{g}})$. Firstly, in this generality it is natural to replace $\chi_{q}(L({\boldsymbol{\psi}}))$ with $\chi_{q}(L^{\neq 0}({\boldsymbol{\psi}}))$ in the left-hand side (these two quantities are equal for highest $\ell$-weights for which the action extends to the entire $U_{q}(L{\mathfrak{g}})$). Secondly, it is naive to hope that the right-hand side should be something as nice as the Euler characteristic. Therefore, our intended generalization will be to find a geometric formula for the coefficients of $\chi_{q}(L^{\neq 0}({\boldsymbol{\psi}}))$, in terms of the algebraic variety $N^{\text{stab}}_{\boldsymbol{x},{\boldsymbol{m}},{\mathbf{l}}}$. Since we already know from (6) that the aforementioned coefficients are given by the dimensions of the vector spaces ${\mathcal{S}}_{{\boldsymbol{n}}}/({\mathcal{S}}_{{\boldsymbol{n}}}\cap\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}})$, then the problem becomes equivalent to the following:
@@ 1.5. General highest $\ell$-weights $\boldsymbol{\psi}$}
To set up our solution to the problem above, in what follows we fix arbitrary
@@
For any symmetrizable Kac-Moody $\mathfrak{g}$}, we have a ring isomorphism
@@
for any ${\boldsymbol{n}},\boldsymbol{x},{\boldsymbol{m}}$ as above, where ${\boldsymbol{\psi}}^{\emph{den}}$ denotes the denominator of (7). The space $M_{\boldsymbol{x},{\boldsymbol{m}}}^{\emph{stab}}$ has no odd $K$-theory.
@@
(15) ${\mathcal{V}}_{{\boldsymbol{n}}}\Big/\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}\xrightarrow{\sim}(\Delta_{{\mathbf{l}}}\cdot{\mathcal{V}}_{{\boldsymbol{n}}})\Big/\left((\Delta_{{\mathbf{l}}}\cdot{\mathcal{V}}_{{\boldsymbol{n}}})\cap\Theta\left(\frac{1}{{\boldsymbol{\psi}}^{\text{den}}}\right)_{\boldsymbol{x}}\right)\subseteq{\mathcal{V}}_{{\boldsymbol{n}}}\Big/\Theta\left(\frac{1}{{\boldsymbol{\psi}}^{\text{den}}}\right)_{\boldsymbol{x}}$
@@
For example, it is quite conceivable that for the reachable highest $\ell$-weights $\boldsymbol{\psi}$} to which Theorem 1.4 applies, the direct image map
@@
is injective and its image coincides with the vector space that appears on the first line of (17). If $N_{\boldsymbol{x},{\boldsymbol{m}},{\mathbf{l}}}^{\text{stab}}$ could also be shown to not have any odd $K$-theory, then the previous sentence would provide an alternate proof of Theorem 1.4.
@@ 1.7. Critical $K$-theory
Let us consider critical $K$-theory, which was developed in [1, 12, 28] (and more recently [6]) and applied to the understanding of certain simple modules $L({\boldsymbol{\psi}})$ in [47, 48]; see also [43] for the original definition of $K$-theoretic Hall algebras. We will start from a certain quiver with potential which has been studied in [22]: let $Q$ have vertex set $I$ and a single arrow $_{j}\square_{i}:i\rightarrow j$ for all $i,j\in I$ (this includes a single loop at every vertex $i$). Then we consider the potential
@@
where the spaces in the bottom row are analogues of those in the top row for the quiver of Theorem 1.4, intersected with the stable open locus (see Subsection 4.6 for more details). In the bottom row, $\Delta_{{\mathbf{l}}}^{\text{gr}}$ is (a slightly renormalized version of) the product of those linear factors of (14) which vanish at $z_{ia}=x_{ia}$, specifically
@@
For any strongly symmetrizable  [^2] See (65) for the definition of strongly symmetrizable Kac-Moody Lie algebras. We note that all finite type and affine type Lie algebras are strongly symmetrizable, except for $\widehat{{\mathfrak{sl}}}_{2}$. Kac-Moody Lie algebra $\mathfrak{g}$} and any highest $\ell$-weight $\boldsymbol{\psi}$}, the image of the composition $\rightarrow\rightarrow\downarrow$ in diagram (20) is naturally isomorphic to the vector space
@@
We conjectured the following result in the first draft of the present paper, and were later informed by Cao-Okounkov-Zhou-Zhou that it follows from their general theory of stable envelopes in critical $K$-theory ([5]).
Theorem 1.9.
(conjectured by the author, proved by [5]) The maps $\pi$ and $\iota^{\emph{gr},\emph{stab}}_{\boldsymbol{x},{\boldsymbol{m}}}$ in (20) are surjective and injective, respectively.
Combining Theorems 1.8 and 1.9 with statement (22) implies the following.
Corollary 1.10.
For strongly symmetrizable $\mathfrak{g}$} and any highest $\ell$-weight $\boldsymbol{\psi}$}, we have
(24) ${\mathcal{S}}_{{\boldsymbol{n}}}\Big/\left({\mathcal{S}}_{{\boldsymbol{n}}}\cap\Theta({\boldsymbol{\psi}})_{\boldsymbol{x}}\right)\cong\left[\Delta^{\emph{gr}}_{{\mathbf{l}}}\cdot K(Q^{\emph{gr}},W^{\emph{gr}})^{\emph{stab}}_{{\boldsymbol{m}},\boldsymbol{x}}\right]\otimes_{{\mathbb{Q}}}{\mathbb{K}}$
We claim that the isomorphism (24) provides a solution to problems (12) and (17), and thus generalizes and categorifies Theorem 1.4 (using the fact that critical $K$-theory is categorified by derived factorization categories). To see this, note that the critical $K$-theory in the right-hand side of (24) is a replacement of the $K$-theory of the critical locus of $W^{\text{gr}}$, while multiplication by $\Delta^{\text{gr}}_{{\mathbf{l}}}$ is a derived version of imposing the vanishing of the $l_{i,s}$ linear combinations of framed paths that cut out the closed subvariety (9). Thus, the right-hand side of (24) is a replacement of $K(N_{\boldsymbol{x},{\boldsymbol{m}},{\mathbf{l}}}^{\text{stab}})\otimes_{{\mathbb{Q}}}{\mathbb{K}}$, which in turn categorifies the Euler characteristic in (8). Meanwhile, the left-hand side of (24) categorifies the coefficients of the $q$-character of $L({\boldsymbol{\psi}})$, due to (5).
1.11. Conclusion
To summarize, we provide a geometric construction of all simple modules $L({\boldsymbol{\psi}})$ in the Borel category $\mathcal{O}$} of quantum loop algebras $U_{q}(L{\mathfrak{g}})$, for any strongly symmetrizable Kac-Moody Lie algebra $\mathfrak{g}$} (which includes all finite and affine $\mathfrak{g}$} except for $\widehat{{\mathfrak{sl}}}_{2}$). This fits into a very active field of research, that started with the Euler characteristic construction of [26], continued with the connections with physics explored in [4], and more recently involved the construction of Kirillov-Reshetikhin and prefundamental modules via critical $K$-theory in [47, 48]. While we consider similar objects as the latter two papers, our approach is somewhat different, and in a certain sense complementary to theirs. To explain this last comment, recall that in loc. cit. the authors defined a convolution product on
@@ 1.11. Conclusion
related it to the quantum affine algebra $U_{q}(L{\mathfrak{g}})$, and defined prefundamental modules (i.e. the case ${\mathbf{l}}={\boldsymbol{0}}$ in (7)) using critical $K$-theory for the same quiver with potential as us. On the other hand, our viewpoint is to start from the vector space
@@
which we already know from [38] yields the simple module $L({\boldsymbol{\psi}})$, and then seek geometric realizations for the direct summands in (26). With this in mind, our Corollary 1.10 not only provides a geometric construction of all simple modules of quantum affine algebras, but provides an explicit computation of the corresponding critical $K$-theory in terms of shuffle algebras. We note that Proposition 4.2 also provides a shuffle algebra computation of the critical K-HA studied in [48].
@@ 1.12. Acknowledgements
I would like to thank Ben Davison, David Hernandez, Shivang Jindal, Hiraku Nakajima, Tudor Pădurariu, Markus Reineke and Yukinobu Toda for numerous important conversations on the topics of the present paper, and the authors of [5] for their proof of Theorem 1.9.
@@ 2. Moduli of quiver representations
We recall the cell decomposition of moduli spaces of stable framed quiver representations from [13], and use it to describe the ring structure on the $K$-theory of the aforementioned moduli spaces (in a different way from the earlier work [16]). Hereafter, we write $K(X)$ for the 0-th $K$-theory  [^3] The contents of the present Section hold equally well for topological and algebraic $K$-theory; the former is closer in spirit to [26], while the latter is closer technically to Section 4. ring of a smooth complex algebraic variety $X$, with rational coefficients. We thank Markus Reineke for sharing with us much advice pertaining to the contents of the present Section.